OpenStudy (moongazer):

A temple was built in 50 BC and burnt down 75 years later. In what year did it burn down?

OpenStudy (dumbcow):

25 AD

OpenStudy (moongazer):

nope

OpenStudy (moongazer):

26AD

OpenStudy (moongazer):

26 AD. There was no year 0. that is the answer according to the online quiz that i answered

OpenStudy (dumbcow):

haha, nice there was no year 0 AD

OpenStudy (anonymous):

ah I see

OpenStudy (moongazer):

but i can't still understand why

OpenStudy (anonymous):

i smelt the trap, so waited for another to fall in it X)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

virtual nose

OpenStudy (moongazer):

please explain further why it is the answer ^_^

OpenStudy (dumbcow):

BC is before Christ AD is a latin phrase which means In the year of our Lord well when Christ was born, it began the 1st " year of our Lord" year 1, thus no year 0

OpenStudy (anonymous):

-50 -49 ... -1, 1, ... , 26. ^ first year you've to count so it's 49 + 26 = 75 years after

OpenStudy (dumbcow):

formula?? i dont know, add 1 i guess

OpenStudy (moongazer):

no, i am referring to the solution

OpenStudy (moongazer):

sorry

OpenStudy (dumbcow):

well like Reku said 75 - 50+1 = 26 If the BC year is B, AD year is A, and n is num of years between -> A = (n - B+1)